# Dirichlet moment tensors and the correspondence between admixture and mixture of product models

### v1

@@ 2.3 Correspondence between spaces of LDA and mixture models
    For all $N$ and $\overline{\alpha}$, an upper bound for $C_{1}(N,\overline{\alpha}$ is readily available:
@@
    By noticing that $C_{1}(N,\overline{\alpha})$ is the expected number of sets in the random partition of $\{1,\dots,N\}$ with the Ewens distribution, we can explicitly calculate (see, e.g., [20])
    $C_{1}(N,\overline{\alpha})=\dfrac{\overline{\alpha}}{\overline{\alpha}}+\dfrac{\overline{\alpha}}{\overline{\alpha}+1}+\dots+\dfrac{\overline{\alpha}}{\overline{\alpha}+N-1}\asymp\overline{\alpha}\log(N)$
    as $N\rightarrow\infty$.
@@
    This proposition will prove useful for establishing the identifiability of the LDA model (cf. Corrolary 1), and moreover for investigating its parameter contraction behavior using the following approach. First, the posterior contraction rates for the density functions may be obtained using standard techniques (cf. [29]), where the main requirements are an upper bound on the covering number of the space of LDA models and a lower bound of the KL ball around the true model. Both can be provided by combining (16) with the existing bounds for mixtures of products [52, 69] (see Section 4). Second, to transfer the density contraction rate to that of the parameter, we are going to lower bound the distance between LDA’s densities by the distance between parameters. This step is considerably challenging, but thanks to (17), it can be done by lower bounding the distance between mixtures of products’ densities instead, which is somewhat more amenable. A collection of those inverse bounds in different settings will be presented in Section 3.
@@
    It is natural to ask whether the multiplicative constants in these inequalities make the posterior contraction rates of the LDA model worse than the mixture of products, especially in the regime where the document length $N$ is large. We highlight that in the first step of showing density contraction rate, the multiplicative constant in inequality (16) only affects the covering number of the space of LDA models $C_{1}(N,\overline{\alpha})$ times larger than the space of mixtures of products, which is $C_{1}(N,\overline{\alpha})$ times larger than the space of parameters (see, e.g. [52, 69]). Hence, the Le Cam dimension of the model space [44, 65], which is practically the logarithm of the covering number, only worsens by $\log(C_{1}(N,\overline{\alpha}))$, which only affects the density contraction rate by that same amount. In the second step of lower bounding the distance between densities, it suffices to use the constant $C_{1}(N,\overline{\alpha})$ for the minimally admissible $N$ for which the inverse bounds of mixtures of products can be obtained. Such $N$ can be as small as 2K12K-1 in the general setting and as 3 or 4 in the setting where topics are linearly independent.
@@ 3.2 Finer identifiability via inverse bounds
for some $r\geq 1$, and the multiplicative constant in this inequality depends only on $G_{0}$ and $K$. The name inverse bounds come from the fact that bounds in the opposite direction of (25) (i.e., upper bound distance between densities by distance between parameters; Appendix D.1 and D.2) are relatively easier to obtain. It can be seen that these inverse bounds are stronger than identifiability because the LHS of the inequality (25) is zero implies the RHS is zero. Thus, inverse bounds represent a finer characterization of identifiability.
@@ 4.1 Posterior contraction of densities and parameters
The proof of this theorem appears in Appendix D.3. We add some remarks.
@@ Acknowledgement
We are grateful to Dr. Yun Wei (UT Dallas) for constructive comments that helped to improve the presentation of the paper and to Dr. Viet Cuong Do (VNU) for helpful algebraic references in the early stage of the project. This research was supported in part by the National Science Foundation, via grants DMS-2015361 (XLN), DMS-2052653 (JT), the National Institute of General Medical Sciences of the NIH under award number R35GM151145, and a research gift from Wells Fargo (XLN). The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
@@
Outline of the appendix: The appendix is organized as follows. In Appendix A, we provide the proof of the first part of Lemma 1 regarding representation of the Dirichlet moment tensors by diagonal tensors (by showing Equation (31)). In Appendix B, we prove Proposition 2 (recursive computation of the Dirichlet moment tensors) and discuss the connection of this proposition with the tensor representation result from Lemma 1. Appendix C provides the proof of Theorem 2, characterizing the minimal document length required for inverse bounds under exact and over-fitted scenarios under different assumptions on the true topics. In Appendix D, we provide the proof of the posterior contraction results for density and parameter estimation, in particular Theorem 3. This section also contains the relevant results on the geometry of multinomial models, providing upper bounds for various metrics on the space of multinomial densities. Appendix E provides the proof for the posterior contraction of the latent allocation, in particular, Theorem 4. Finally, Appendix F contains proofs of other results, including Proposition 3 on minimal document lengths for identifiability, and Lemma 5, establishing properties on various metrics on the space of distributions, which is used in the proof of Proposition 1.
@@ A.1 Proof of Equation (11)
Appendix B Proof and discussion of Proposition 2
B.1 Proof of Proposition 2
@@
for any $x\in\mathbb{R}^{K}$ and permutation $T_{(S_{1},\dots,S_{\ell})}$ such that $|S_{i}|=d_{i}$. Now, we encode each partition by frequency instead of composition, i.e., we correspond each tuple $(d_{1},\dots,d_{\ell})$ to tuple $(p_{1},\dots,p_{N})$ such that $p_{j}=\#\{i:d_{i}=j\}$. Acting $x,\dots,x]$ to both sides of Eq. (B.1) yields
@@
B.2 Discussion on the connection between Proposition 2 and Lemma 1
Having seen that Lemma 1 implies Proposition 2, we wrap up this section by talking about the inverse implication. That is, we are going to show that if Proposition 2 holds, then the Dirichlet moment tensor $Q^{(N)}_{\alpha}$ and the tensor
@@ B.2 Discussion on the connection between Proposition 2 and Lemma 1
which uniquely determines the polynomial $Q_{N}(x)$ from $\{Q_{n}(x):n\leq N-1\}$. Proposition 2 states that $Q^{(N)}_{\alpha}[x,\dots,x]$ is a polynomial that satisfies this recursive formula. However, in the proof of Proposition 2 after the first step, we only make use of some algebraic transformations of the RHS of Eq. (B.1) to show that $\tilde{Q}^{(N)}_{\alpha}[x,\dots,x]$ also satisfies this recursive formula (without using Lemma 1). Hence
@@
Appendix C Proof of Theorem 2
@@ C.1 Proof of Theorem 2(a)
C.1.1 Proof of the exact-fitted inverse bound for distinct topics
@@
C.1.2 Proof of the over-fitted inverse bound for distinct topics
@@
C.2 Proof of Theorem 2 (b): Linearly independent topics
@@
C.2.1 Proof of the exact-fitted inverse bound for linearly independent topics
@@
C.2.2 Proof of the over-fitted inverse bound for linearly independent topics
@@
C.3 Proof of Theorem 2 (c): With the anchor-word assumption
@@
C.3.1 Proof of the exact-fitted inverse bound for anchor-word topics
@@
Appendix D Proofs of convergence for densities and parameter estimation
@@
D.1 Geometry of Multinomial models
@@ 1. For any two probability vectors $\theta,\theta^{\prime}\in\Psi$ and positive integer $N$, we have:
D.2 Relevant Results for the Posterior Contraction Theorem
@@ D.3 Proof of Theorem 3
Appendix E Proof of Theorem 4
E.1 Setup and organization of the proof of Theorem 4
@@
E.2 Posterior contraction rate for density estimation
@@
where the last inequality uses the type-II error control for the tests $\Phi_{mN}^{1}$ and $\Phi_{\tilde{N}}^{2}$. Hence, the expectation of the third term in the right side of the decomposition in Eq. (E.2) is upper bounded by
@@
E.3 Posterior contraction rate for allocation $q$
@@
Appendix F Proof of other theoretical results
F.1 Proof of Proposition 3
@@
F.2 On the distances between distributions

### v3

@@ 2.3 Correspondence between spaces of LDA and mixture models
    For all $N$ and $\overline{\alpha}$, an upper bound for $C_{1}(N,\overline{\alpha})$ is readily available:
@@
    Meanwhile, the constant $C_{2}(N,\overline{\alpha})$ involves a sum over all partitions of $[N]$ and is harder to bound by a simple expression of $N$ and $\overline{\alpha}$. To understand the tightness of those bounds of $C_{1}$ and $C_{2}$, we have conducted simulation studies in Appendix B for various settings of $N$, $\overline{\alpha}$ and $K$. When $\overline{\alpha}<1$, we find that $C_{1}$ and $C_{2}$ are both close to 1, which indicates that tighter theoretical bounds might be obtained.
@@
    This proposition will prove useful for establishing the identifiability of the LDA model (cf. Corollary 1), and moreover for investigating its parameter contraction behavior using the following approach. First, the posterior contraction rates for the density functions may be obtained using standard techniques (cf. [29]), where the main requirements are an upper bound on the covering number of the space of LDA models and a lower bound of the KL ball around the true model. Both can be provided by combining (16) with the existing bounds for mixtures of products [52, 69] (see Section 4). Second, to transfer the density contraction rate to that of the parameter, we are going to lower bound the distance between LDA’s densities by the distance between parameters. This step is considerably challenging, but thanks to (17), it can be done by lower bounding the distance between mixtures of products’ densities instead, which is somewhat more amenable. A collection of those inverse bounds in different settings will be presented in Section 3.
@@
    It is natural to ask whether the multiplicative constants in these inequalities make the posterior contraction rates of the LDA model worse than the mixture of products, especially in the regime where the document length $N$ is large. We highlight that in the first step of showing density contraction rate, the multiplicative constant in inequality (16) only affects the covering number of the space of LDA models $C_{1}(N,\overline{\alpha})$ times larger than the space of mixtures of products, which is $C_{1}(N,\overline{\alpha})$ times larger than the space of parameters (see, e.g. [52, 69]). Hence, the Le Cam dimension of the model space [44, 65], which is practically the logarithm of the covering number, only worsens by $\log(C_{1}(N,\overline{\alpha}))$, which only affects the density contraction rate by that same amount. In the second step of lower bounding the distance between densities, it suffices to use the constant $C_{2}(N,\overline{\alpha})$ for the minimally admissible $N$ for which the inverse bounds of mixtures of products can be obtained. Such $N$ can be as small as 2K12K-1 in the general setting and as 3 or 4 in the setting where topics are linearly independent.
@@ 3.2 Finer identifiability via inverse bounds
for some $r\geq 1$, and the multiplicative constant in this inequality depends only on $G_{0}$ and $K$. The name inverse bounds come from the fact that bounds in the opposite direction of (25) (i.e., upper bound distance between densities by distance between parameters; Appendix E.1 and E.2) are relatively easier to obtain. It can be seen that these inverse bounds are stronger than identifiability because the LHS of the inequality (25) is zero implies the RHS is zero. Thus, inverse bounds represent a finer characterization of identifiability.
@@ 4.1 Posterior contraction of densities and parameters
The proof of this theorem appears in Appendix E.3. We add some remarks.
@@ Acknowledgement
We are grateful to Dr. Yun Wei (University of Texas at Dallas) for many constructive comments and discussions that helped improve the paper, and for pointing out many typos. Dat Do is grateful to Dr. Yun Wei for discussing Lemma A.8 in [68] (arXiv v1) and bringing up reference [66]. We thank Dr. Viet Cuong Do (Vietnam National University) for helpful discussions on algebraic geometry and for providing us with many important algebra references. This research was supported in part by the National Science Foundation, via grants DMS-2015361 (XLN), DMS-2052653 (JT), the National Institute of General Medical Sciences of the NIH under award number R35GM151145, and a research gift from Wells Fargo (XLN). The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
@@
Outline of the appendix: The appendix is organized as follows. In Appendix A, we provide the proof of the first part of Lemma 1 regarding representation of the Dirichlet moment tensors by diagonal tensors (by showing Equation (31)). In Appendix B, we examine the magnitude of constants $C_{1}$ and $C_{2}$ in Proposition 1 by simulations. In Appendix C, we prove Proposition 2 (recursive computation of the Dirichlet moment tensors) and discuss the connection of this proposition with the tensor representation result from Lemma 1. Appendix D provides the proof of Theorem 2, characterizing the minimal document length required for inverse bounds under exact and over-fitted scenarios under different assumptions on the true topics. In Appendix E, we provide the proof of the posterior contraction results for density and parameter estimation, in particular Theorem 3. This section also contains the relevant results on the geometry of multinomial models, providing upper bounds for various metrics on the space of multinomial densities. Appendix F provides the proof for the posterior contraction of the latent allocation, in particular, Theorem 4. Finally, Appendix G contains proofs of other results, including Proposition 3 on minimal document lengths for identifiability, and Lemma 5, establishing properties on various metrics on the space of distributions, which is used in the proof of Proposition 1.
@@ A.1 Proof of Equation (11)
Appendix B Discussions and Numerical Experiments for the constants in Proposition 1
We conduct a simple simulation study to investigate the constants in Proposition 1. In particular, we want to understand the magnitude of $C_{1}$ and $C_{2}$ in the inequalities:
$d_{TV}(p^{\mathcal{L}}_{\alpha,\Theta},p^{\mathcal{L}}_{\alpha^{\prime},\Theta^{\prime}})\leq C_{1}(N,\overline{\alpha})\times d_{TV}(p^{\mathcal{M}}_{\tilde{\alpha},\Theta},p^{\mathcal{M}}_{\tilde{\alpha}^{\prime},\Theta^{\prime}})$ (47)
and
$d_{TV}(p^{\mathcal{M}}_{\tilde{\alpha},\Theta},p^{\mathcal{M}}_{\tilde{\alpha}^{\prime},\Theta^{\prime}})\leq C_{2}(N,\overline{\alpha})\times d_{TV}(p^{\mathcal{L}}_{\alpha,\Theta},p^{\mathcal{L}}_{\alpha^{\prime},\Theta^{\prime}}),$ (48)
for every pair of $(\alpha,\Theta)$ and $(\alpha^{\prime},\Theta^{\prime})$.
In this simulation study, we set $K=3,V=10,N=5$, and randomly generate pairs $\tilde{\alpha},\tilde{\alpha}^{\prime}\sim\mathrm{Dir}_{(1,1,1)}$ (uniform on the simplex $\Delta^{K-1}$) and all topics $\Theta_{k},\Theta^{\prime}_{k}\sim\mathrm{Dir}_{(1,\dots,1)}$ (uniform on the simplex $\Delta^{V-1}$) for $k=1,\dots,K$. We then construct $p^{\mathcal{M}}$ and $p^{\mathcal{L}}$ as $V\times V\dots\times V$ (NN times) tensor and compute $d_{TV}(p^{\mathcal{M}}_{\tilde{\alpha},\Theta},p^{\mathcal{M}}_{\tilde{\alpha}^{\prime},\Theta^{\prime}})$ as half of the $\ell_{1}$ distance between two tensors (similar for $d_{TV}(p^{\mathcal{L}}_{{\alpha},\Theta},p^{\mathcal{L}}_{{\alpha}^{\prime},\Theta^{\prime}})$). We generate 1000 pairs of parameters and plot them in Figure 5 with the corresponding estimated lower and upper slope $C_{1}$ and $C_{2}$. We consider two settings where $\overline{\alpha}=1$ and $\overline{\alpha}=0.1$.
(a) Setting $N=5,K=3$, and $\overline{\alpha}=1$
(a) *
(b) Setting $N=5,K=3$, and $\overline{\alpha}=0.1$
(b) *
Figure 5: Comparing $d_{TV}$ between Mixture models and corresponding Admixture models.
It can be seen that the estimated $C_{1}$, which is calculated by the minimum of $\dfrac{d_{TV}(p^{\mathcal{L}}_{{\alpha},\Theta},p^{\mathcal{L}}_{{\alpha}^{\prime},\Theta^{\prime}})}{d_{TV}(p^{\mathcal{M}}_{\tilde{\alpha},\Theta},p^{\mathcal{M}}_{\tilde{\alpha}^{\prime},\Theta^{\prime}})}$ over all pairs of $(\alpha,\Theta)$ and $(\alpha^{\prime},\Theta^{\prime})$ is very close to 1. On the other hand, $C_{2}$, which is calculated by the minimum of $\dfrac{d_{TV}(p^{\mathcal{M}}_{\tilde{\alpha},\Theta},p^{\mathcal{M}}_{\tilde{\alpha}^{\prime},\Theta^{\prime}})}{d_{TV}(p^{\mathcal{L}}_{{\alpha},\Theta},p^{\mathcal{L}}_{{\alpha}^{\prime},\Theta^{\prime}})}$ over all pairs of $(\alpha,\Theta)$ and $(\alpha^{\prime},\Theta^{\prime})$ is around $\dfrac{3}{2}$ when $\overline{\alpha}=1$ and is close to 1 when $\overline{\alpha}=0.1$ ($1/C_{2}$ is displayed in Figure 5 because $d_{TV}(p^{\mathcal{L}}_{{\alpha},\Theta},p^{\mathcal{L}}_{{\alpha}^{\prime},\Theta^{\prime}})$ is plotted in the x-axis and $d_{TV}(p^{\mathcal{M}}_{\tilde{\alpha},\Theta},p^{\mathcal{M}}_{\tilde{\alpha}^{\prime},\Theta^{\prime}})$ is plotted in the y-axis). Those estimates together show that the geometry of the mixture models and admixture models are quite close, and a better bound for the constants $C_{1}$ and $C_{2}$ could be obtained with more sophisticated mathematical techniques. Finally, it is worth emphasizing that the constants $C_{1}$ and $C_{2}$ should not be confused with the constant $C$ appearing in the parameter estimation rates of both models in Theorem 3, which only depends on the smallest $N$ so that the inverse bounds (Theorem 2) hold – this can be as small as 3 (or 4) in the exact-fitted (or over-fitted) setting and linear independent topics (see Remark 6).
Appendix C Proof and discussion of Proposition 2
C.1 Proof of Proposition 2
@@
for any $x\in\mathbb{R}^{K}$ and permutation $T_{(S_{1},\dots,S_{\ell})}$ such that $|S_{i}|=d_{i}$. Now, we encode each partition by frequency instead of composition, i.e., we correspond each tuple $(d_{1},\dots,d_{\ell})$ to tuple $(p_{1},\dots,p_{N})$ such that $p_{j}=\#\{i:d_{i}=j\}$. Acting $x,\dots,x]$ to both sides of Eq. (C.1) yields
@@
C.2 Discussion on the connection between Proposition 2 and Lemma 1
Having seen that Lemma 1 implies Proposition 2, we now talk about the inverse implication. That is, we are going to show that if Proposition 2 holds, then the Dirichlet moment tensor $Q^{(N)}_{\alpha}$ and the tensor
@@ C.2 Discussion on the connection between Proposition 2 and Lemma 1
which uniquely determines the polynomial $Q_{N}(x)$ from $\{Q_{n}(x):n\leq N-1\}$. Proposition 2 states that $Q^{(N)}_{\alpha}[x,\dots,x]$ is a polynomial that satisfies this recursive formula. However, in the proof of Proposition 2 after the first step, we only make use of some algebraic transformations of the RHS of Eq. (C.1) to show that $\tilde{Q}^{(N)}_{\alpha}[x,\dots,x]$ also satisfies this recursive formula (without using Lemma 1). Hence
@@
Appendix D Proof of Theorem 2
@@ D.1 Proof of Theorem 2(a)
D.1.1 Proof of the exact-fitted inverse bound for distinct topics
@@
D.1.2 Proof of the over-fitted inverse bound for distinct topics
@@
D.2 Proof of Theorem 2 (b): Linearly independent topics
@@
D.2.1 Proof of the exact-fitted inverse bound for linearly independent topics
@@
D.2.2 Proof of the over-fitted inverse bound for linearly independent topics
@@
D.3 Proof of Theorem 2 (c): With the anchor-word assumption
@@
D.3.1 Proof of the exact-fitted inverse bound for anchor-word topics
@@
Appendix E Proofs of convergence for densities and parameter estimation
@@
E.1 Geometry of Multinomial models
@@ 1. For any two probability vectors $\theta,\theta^{\prime}\in\Psi$ and positive integer $N$, we have:
E.2 Relevant Results for the Posterior Contraction Theorem
@@ E.3 Proof of Theorem 3
Appendix F Proof of Theorem 4
F.1 Setup and organization of the proof of Theorem 4
@@
F.2 Posterior contraction rate for density estimation
@@
where the last inequality uses the type-II error control for the tests $\Phi_{mN}^{1}$ and $\Phi_{\tilde{N}}^{2}$. Hence, the expectation of the third term in the right side of the decomposition in Eq. (F.2) is upper bounded by
@@
F.3 Posterior contraction rate for allocation $q$
@@
Appendix G Proof of other theoretical results
G.1 Proof of Proposition 3
@@
G.2 On the distances between distributions